If a polynomial p(x) divided by x-2 then remainder is one and when it is divided by x+1 the remainder is -2 find the remainder when the given polynomial divided by (x-2)(x+1)


1
Expert's answer
2021-05-18T15:04:19-0400

Note that "(x-2)(x+1)=x^2-x-2" is quadratic, so if we divide another polynomial by it then the remainder will be 0, a non-zero constant or a linear polynomial. Any remainder of greater degree could be divided further.

Suppose "p(x)=(x-2)(x+1)g(x)+kx+b"

Then


"\dfrac{p(x)}{x-2}=(x+1)g(x)+\dfrac{kx+b}{x-2}"

"=(x+1)g(x)+k+\dfrac{2k+b}{x-2}"


If a polynomial "p(x)" is divided by x-2 then remainder is one.

Hence "2k+b=1"


"\dfrac{p(x)}{x+1}=(x-2)g(x)+\dfrac{kx+b}{x+1}"

"=(x-2)g(x)+k+\dfrac{-k+b}{x+1}"


If a polynomial "p(x)" is divided by x+1 then remainder is "-2."

Hence "-k+b=-2."


"\begin{matrix}
 2k+b=1 \\
 -k+b=-2
\end{matrix}"

"\begin{matrix}
 3k=3 \\
 b=k-2
\end{matrix}"

"\begin{matrix}
 k=1 \\
 b=-1
\end{matrix}"

So the remainder when the given polynomial divided by (x-2)(x+1) is "x-1."